OpenStudy (anonymous):

describe a transformation that changes f(x) to G(x) 1. f(x)=x+3 g(x)=x-2 2.f(x)=-x-4 g(x)=x+1

OpenStudy (anonymous):

g(x)=x-2=x+3-3-2=x+3-5=f(x)-5

OpenStudy (anonymous):

there is two separate questions which is which

OpenStudy (anonymous):

first

OpenStudy (anonymous):

2.g(x)=x+1=-(-x)+1=-(-x-4+4)+1=-(-x-4)-4+1=-(-x-4)-3=-f(x)-3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw